OpenStudy (anonymous):

I just need some guidance on how to get the answer! please. :) Solve for x. 8x - 4 = 2x - 3

OpenStudy (callisto):

First, subtract 2x from both sides, what do you get?

OpenStudy (anonymous):

6x - 4 = -3?

OpenStudy (callisto):

Yes. Now, add 4 to both sides, what do you get?

OpenStudy (anonymous):

6x - 8 = 1?

OpenStudy (valpey):

try -4+4

OpenStudy (callisto):

Nope. Add 4 to both sides => 6x - 4 +4 = -3 +4 => ... ?

OpenStudy (anonymous):

Ha sorry im a blonde! :P hold on let me try it

OpenStudy (anonymous):

i got 6x to 1

OpenStudy (anonymous):



OpenStudy (callisto):

Yes :) Now, divide both sides by 6. What do you get?

OpenStudy (anonymous):

i got 6 haha and 1

OpenStudy (callisto):

Hmm???

OpenStudy (anonymous):

1 = 6

OpenStudy (anonymous):

i mean x=6 haha sory

OpenStudy (callisto):

Nope... Not x =6...